# Fast Penalized Generalized Estimating Equations for Large Longitudinal Functional Datasets

### v1

@@ 2.3 Working Correlations
If $\mathbb{V}_{i}(s)$ has the AR1 structure $\text{Cor}(Y_{i,j}(s),Y_{i,j^{\prime}}(s)\mid\mathbf{X}_{i,j})=\rho(s)^{|j-j^{\prime}|}$, $\mathbb{V}_{i}^{-1}(s)$ can be efficiently computed because, for $\rho(s)\geq 0$, its decomposition yields a Toeplitz $\mathbf{R}_{i}(s)$. For an estimated $\widehat{\bm{r}}_{i}(s)=[\bm{Y}_{i}(s)-\widehat{\bm{\mu}}_{i}(s)]\widehat{\mathbf{A}}_{i}(s)^{-1/2}$, we can quickly calculate $\widehat{\mathbb{V}}_{i}(s)^{-1}\widehat{\bm{r}}_{i}(s)$ by solving the Toeplitz system $\bm{a}=\widehat{\mathbf{R}}_{i}(s)\widehat{\bm{r}}_{i}(s)$ with, for example, the generalized Schur algorithm (Ammar and Gragg,, 1988). This is done without fully constructing the $n_{i}\times n_{i}$ matrix $\widehat{\mathbb{V}}_{i}(s)$. When data are observed irregularly, one can use the algorithm proposed in Allévius, (2018).
Correlation parameters
@@
When adopting the block exchangeable or AR1 correlation structures, each ${\mathbb{V}}_{i}(s)$ is a function of a nuisance correlation parameter $\rho(s)$. We estimate these at each point $s$ separately and then optionally smooth over the functional domain to reduce variability. Defining residuals as $e_{i,j}(s)=\frac{Y_{i,j}(s)-{\mu}_{i,j}(s)}{\sqrt{v(\mu_{i,j}(s))}}$, we use the method of moments estimator for the exchangeable structure (Molenberghs et al.,, 2005): $\widehat{\rho}(s)=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{n_{i}(n_{i}-1)}\sum_{j\leq k}\hat{e}_{i,j}(s)\hat{e}_{i,k}(s)$, and truncate the $\widehat{\rho}(s)$ at $1-\epsilon$ or $-1+\epsilon$ if they fall outside the $(-1,1)$ range. For an AR1 structure, we estimate each $\rho_{i}(s)$ with the Yule-Walker equations (Yule,, 1927; Walker,, 1931) when the longitudinal observations are sampled at regular time intervals. If sampled irregularly, we estimate $\rho_{i}(s)$ with the MLE estimator proposed in Allévius, (2018). We then calculate $\widehat{\rho}(s)=\frac{1}{N}\sum_{i=1}^{N}\widehat{\rho}_{i}(s)$ and truncate the $\widehat{\rho}(s)$ at 00 or $1-\epsilon$ if they fall outside the $[0,1)$ range. The $\widehat{\rho}(s)$ are calculated twice in our framework: 1) first using $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ to calculate $\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})$ that is plugged into the one-step estimator (3), and 2) second using $\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}$ to calculate $\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)})$ that is plugged into the $\widehat{\text{Var}}\left(\widehat{\bm{\beta}}_{\Lambda_{1}}^{(1)}\right)$ estimator (see expression (6)).
2.4 Tuning $\Lambda$
@@
To calculate an initial estimate of $\widehat{\bm{\theta}}^{(0)}_{\Lambda_{0}}$, we select the smoothing parameters, denoted as $\Lambda_{0}$, with fast restricted maximum likelihood (Wood,, 2011). We found, however, that calculating the one-step estimate with the same $\Lambda_{0}$ values (i.e. $\bm{\theta}^{(1)}_{\Lambda_{0}}$) tends to produce inaccurate coefficient estimates. Therefore, we propose to tune the smoothing parameters for the one-step, denoted as $\Lambda_{1}$, based on the cross-validated prediction performance of the one-step estimator. In our simulations, K-fold cluster cross-validation (CV) produced one-step estimates with better estimation accuracy than one-step estimates based on smoothing parameters selected with GCV [^2] Restricted Maximum Likelihood cannot be used to select smoothing parameters for fGEE because there is no likelihood to maximize (Chen et al.,, 2013). or the bootstrap-based procedure proposed in Chen et al., (2013). We use the negative log-likelihood as a CV fit criteria. We propose the following scalable CV for large datasets.

### v2

@@ A.3 Efficient inversion of working covariance matrices
If $\mathbb{V}_{i}(s)$ has, for example, then AR1 structure $\text{Cor}(Y_{i,j}(s),Y_{i,j^{\prime}}(s)\mid\mathbf{X}_{i,j})=\rho(s)^{|j-j^{\prime}|}$, $\mathbb{V}_{i}^{-1}(s)$ can also be efficiently computed because, for $\rho(s)\geq 0$, its decomposition yields a Toeplitz $\mathbf{R}_{i}(s)$. For an estimated $\widehat{\bm{e}}_{i}(s)=[\bm{Y}_{i}(s)-\widehat{\bm{\mu}}_{i}(s)]\widehat{\mathbf{A}}_{i}(s)^{-1/2}$, we can quickly calculate $\widehat{\mathbb{V}}_{i}(s)^{-1}\widehat{\bm{e}}_{i}(s)$ by solving the Toeplitz system $\bm{a}=\widehat{\mathbf{R}}_{i}(s)\widehat{\bm{e}}_{i}(s)$ with, for example, the generalized Schur algorithm (Ammar and Gragg, 1988). This is done without fully constructing the $n_{i}\times n_{i}$ matrix $\widehat{\mathbb{V}}_{i}(s)$ or storing $\widehat{\mathbb{V}}^{-1}_{i}(s)$ in memory. When data are observed irregularly, one can use the algorithm proposed in Allévius (2018).
More generally, if $\mathbb{V}_{i,j}$, $\mathbb{V}_{Lon}$ or $\mathbb{V}_{Fun}$ have a number of different common covariance structures (e.g., an exchangeable, AR1, AR(p), moving average MA(q), Gaussian Process covariances when using Matern/RBF kernels) the respective $\mathbf{R}_{i,j}$, $\mathbf{R}_{Lon}$, $\mathbf{R}_{Fun}$ will have Toeplitz structure and can be inverted with the same algorithms. When analyzing a longitudinal functional outcome observed on a multivariate functional domain (e.g., fMRI), the Kronecker identity approach to efficiently invert working covariance matrices described in Section 2.2 can be applied to efficiently invert covariance matrices of the form $\mathbf{R}_{i}=\mathbf{R}_{1}\otimes\mathbf{R}_{2}\otimes\mathbf{R}_{3}$.
A.4 Working correlation parameters
@@
When adopting the block exchangeable or AR1 correlation structures, each ${\mathbb{V}}_{i}(s)$ is a function of a nuisance correlation parameter $\rho(s)$. We estimate these at each point $s$ separately and then optionally smooth over the functional domain to reduce variability. Defining residuals as $e_{i,j}(s)=\frac{Y_{i,j}(s)-{\mu}_{i,j}(s)}{\sqrt{v(\mu_{i,j}(s))}}$, we use the method of moments estimator for the exchangeable structure (Molenberghs et al., 2005): $\widehat{\rho}(s)=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{n_{i}(n_{i}-1)}\sum_{j\leq k}\hat{e}_{i,j}(s)\hat{e}_{i,k}(s)$, and truncate the $\widehat{\rho}(s)$ at $1-\epsilon$ or $1+\epsilon$ if they fall outside the (1,1)(-1,1) range. For an AR1 structure, we estimate each $\rho_{i}(s)$ with the Yule-Walker equations (Yule, 1927; Walker, 1931) when the longitudinal observations are sampled at regular time intervals. If sampled irregularly, we estimate $\rho_{i}(s)$ with the MLE estimator proposed in Allévius (2018). We then calculate $\widehat{\rho}(s)=\frac{1}{N}\sum_{i=1}^{N}\widehat{\rho}_{i}(s)$ and truncate the $\widehat{\rho}(s)$ at 0 or $1-\epsilon$ if they fall outside the $[0,1)$ range. For Kronecker product working covariances, there is only one correlation parameter for $\Sigma_{Lon}$ and one for $\Sigma_{Fun}$. In these cases, the estimator for $\rho_{Lon}$ and $\rho_{Fun}$ is calculated based on a weighted average of the functional domain-specific estimates. For example, for a method of moments estimator for the exchangeable $\rho_{Lon}$, we estimate $\widehat{\rho}_{Lon}=\frac{\sum_{s\in\mathcal{S}}\sum_{i=1}^{N}\sum_{j\leq k}\hat{e}_{i,j}(s)\hat{e}_{i,k}(s)}{\sum_{i=1}^{N}\sum_{s\in\mathcal{S}}\binom{n_{i}}{2}}$.
The correlation parameters (e.g., $\widehat{\rho}(s)$, $\rho_{Lon}$) are calculated twice in our framework: 1) first using $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ to calculate $\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)})$ that is plugged into the one-step estimator (2.2), and 2) second using $\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)}$ to calculate $\widehat{\mathbb{V}}_{i}(\widehat{\bm{\theta}}_{\Lambda_{1}}^{(1)})$ that is plugged into the $\widehat{\text{Var}}\left(\widehat{\bm{\beta}}_{\Lambda_{1}}^{(1)}\right)$ estimator (see main text Section 2.2).
A.5 Functional Principal Component Analysis
Here we discuss estimation of $\mathbf{R}_{i,Fun}$ using functional principal component analysis (FPCA).
Procedurally, we use the initial fit $\widehat{\bm{\theta}}_{\Lambda_{0}}^{(0)}$ to calculate residuals, as used for estimation of other correlation parameters above. We then define $\widehat{e}_{i,j}(s)=\frac{Y_{i,j}(s)-\widehat{\mu}_{i,j}(s)}{\sqrt{v(\widehat{\mu}_{i,j}(s))}}$ and concatenate these into the vectors
$\widehat{\bm{e}}_{i,j}=[\widehat{e}_{i,j}(s_{1}),\ldots,\widehat{e}_{i,j}(s_{L})]^{T}\in\mathbb{R}^{L}$
for all $i\in[N],j\in[n_{i}]$. We pool these across $i\in[N]$ and $j\in[n_{i}]$, ignoring residual dependence, and fit a functional principal component analysis (FPCA) using refund::fpca.face for dense grids and refund::fpca.sc for sparse grids (Xiao et al., 2013, 2016).
We use this to estimate $\mathbf{R}_{i,Fun}$. For example, we denote $\mathbf{C}\{E(s),E(s^{\prime})\}$ as the mean-zero residual process defined along the functional domain (after standardization). We can then use FPCA to approximate the functional covariance kernel
$\mathbf{C}(s,s^{\prime})=Cov\{E(s),E(s^{\prime})\}$
via the eigen-expansion
$\text{Cov}\{E(s),E(s^{\prime})\}=\sum_{k\geq 1}\lambda_{k}\phi_{k}(s)\phi_{k}(s^{\prime})$
where $\phi_{k}(s)$ and $\lambda_{k}$ are eigenfunction and eigenvalue $k$ at $s$, respectively. After fitting the above FPCA, for the observed functional domain point grid $\mathcal{S}$, we denote the matrix of estimated eigenfunctions as $\widehat{\Phi}\in\mathbb{R}^{L\times K}$ where $K$ is the number of retained eigenfunctions, and column $k$ contains $\widehat{\phi}_{k}$ evaluated on the $L\times 1$ grid $\{s_{l}\}_{l=1}^{L}$. We write $\widehat{\Lambda}=\text{diag}(\widehat{\lambda}_{1},\ldots,\widehat{\lambda}_{K})$. Then we estimate our working covariance as
$\widehat{\mathbf{C}}=\widehat{\Phi}\widehat{\Lambda}\widehat{\Phi}^{T}+\hat{\sigma}^{2}I$
where $\hat{\sigma}^{2}$ is the residual variance. In practice, we select the number of eigenfunctions to retain as the smallest $K$ such that $\frac{\sum_{k=1}^{K}\widehat{\lambda}_{k}}{\sum_{k\geq 1}\widehat{\lambda}_{k}}\geq\texttt{pve}$ for percent variance explained, $\texttt{pve}=0.95$. We then take $\widehat{\mathbf{R}}_{i,Fun}=\widehat{\mathbf{C}}_{i}$, where $\widehat{\mathbf{C}}_{i}$ is constructed on the grid of functional domain points observed for cluster $i$. Inverting $\widehat{\mathbf{C}}_{i}$ can be made more computationally efficient by using a Woodbury inversion formula since $\widehat{\mathbf{C}}=\widehat{\Phi}\widehat{\Lambda}\widehat{\Phi}^{T}+\hat{\sigma}^{2}I$ is defined as a low rank (rank $K$) matrix $\widehat{\Phi}\widehat{\Lambda}\widehat{\Phi}^{T}$, plus a diagonal matrix $\hat{\sigma}^{2}I$.
Future extensions could adopt working covariance matrices that define $\widehat{\mathbf{R}}_{i,Fun}$ with cluster-specific estimates of $\widehat{\mathbf{C}}_{i}$ using longitudinal functional PCA methods (Yao et al., 2005; Greven et al., 2010; Shou et al., 2015; Cui et al., 2023) to account for the clustering (e.g., correlation in the observational direction). If one adopts a block diagonal fGEE working covariance matrix that specifies correlation in the functional direction, $\mathbf{R}_{i}=\text{bdiag}(\widehat{\mathbf{R}}_{i,1},\ldots,\widehat{\mathbf{R}}_{i,n_{i}})$, one could, use longitudinal FPCA estiamtes to construct cluster- and longitudinal observation-specific covariance estimates with blocks $\widehat{\mathbf{R}}_{i,j}=\widehat{\mathbf{C}}_{i,j}$.
A.6 Cross-Validation Procedure
A.6.1 Fast Cluster Cross-Validation
@@
We provide a full description of our fast cluster cross-validation (CV) here. We found in simulations that cluster cross-validation produced one-step estimates with better estimation accuracy than one-step estimates based on smoothing parameters selected with the bootstrap-based procedure proposed in Chen et al. (2013). Moreover, restricted Maximum Likelihood cannot be used to select smoothing parameters for fGEE because there is no likelihood to maximize (Chen et al., 2013). We use the negative log-likelihood as a CV fit criteria. We propose the following scalable (K-fold) cluster CV for large datasets.